# Gravitational Waves Emission in Quadratic Gravity: longitudinal modes, angular momentum emission, and positivity of the radiated power

### v1

@@ I Introduction
Gravitational waves are certainly one of the most disruptive discoveries in astrophysics in the recent years. Although its prediction dates back to 1916 [1], the first evidence of its existence was only glimpsed in the 1970s by Hulse and Taylor [2], who observed that the system PSR 1913+16 lost energy and angular momentum in accordance with the theory of General Relativity (GR). It took another 40 years for the first direct detection of a gravitational wave (GW), due to the merger of two black holes, by the LIGO collaboration in 2015 [3]. This event marked the opening of a new window of observations, creating a new perspective for understanding astrophysical phenomena. Two years later, a major breakthrough was achieved with the observation of the merger of two neutron stars by the LIGO-VIRGO collaboration [4]. This was the first event in which both the GW and its electromagnetic counterpart were observed together, marking the beginning of the era of multimessenger astronomy.
@@
Besides allowing the comprehension of astrophysics phenomena from a broader perspective, GW is also a powerful tool for studying the gravitational interaction itself. Although GR is the standard theory for gravitation, with several successful predictions, it fails to describe galaxy rotation curves and the accelerated expansion of the Universe in the presence of ordinary matter [5, 6, 7, 8]; the consistent description of the early (inflationary) Universe is hardly achieved by GR without the presence of extra degrees of freedom [9, 10, 11, 12, 13, 14, 15]. Also, the non-renormalizability and the presence of singularities [16, 17] do not allow us to have a self-consistent quantum version of GR with the standard (perturbative) prescription of quantization. Due to these limitations, there is almost a consensus on the scientific community that GR should not be taken as the definitive theory of gravitation. This motivated the search for new theories of gravity.
@@
The proposition of new theories of gravitation can be done by a plethora of different approaches. They include the modification of the geometry [18, 19, 20, 21, 22], the modification of the gravitational action [23, 24, 25, 26, 27, 28, 29, 30] (in the Riemann manifold), and many others. Usually, all these proposals share one same concern: one should obtain GR (or GR-equivalent theories) as a limiting case. With this in mind, we can understand that one of the simplest extensions to GR, besides considering the cosmological constant, consists of considering contributions to the gravitational action integral of terms that are quadratic in the curvature (this is, for instance, the case of the Starobinsky model of inflation [9]). The presence of quadratic terms in the action (quadratic gravity), however, brings with them a significant modification of the dynamical equations: the field equations are of fourth order in the derivative of the metric tensor.
@@
Theories with higher order derivatives are quite appealing in the context of quantum gravity, and some of the problems of GR may eventually be worked around – for instance, it was shown by Stelle [17], that a renormalizable theory of gravity can be obtained for the quadratic Lagrangian. Despite being renormalizable, quadratic gravity exhibits Ostrogradski instability, which, when quantized, gives rise to ghost-like fields [31]. The quantization of these fields is a nontrivial task and often leads to issues with unitarity [32]. One way to address this issue is to recognize that ghosts are unstable and, as a result, do not contribute to the asymptotic spectrum of the theory [33, 34]. Alternatively, one could define a nontrivial norm for ghost states to restore their unitarity [35, 36]. While there is currently no consensus in the literature on how to handle the ghosts present in higher-order models, quadratic gravity and its extensions [37, 38, 33, 39] can be treated as effective theories that implement corrections to GR.
@@
The investigation of gravitational waves within the framework of quadratic gravity was particularly analyzed in Ref. [40]. [^1] See also Refs. [41, 42, 43, 44, 45, 46] for related works. In this study, a modification to the Einstein-Hilbert action was proposed, incorporating quadratic contributions from the scalar curvature and the Weyl tensor, i.e., $R^{2}$ and $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$ terms, respectively. The GWs obtained in this work exhibited two spin-2 modes—one massive and one massless—along with a scalar (spin-0) mode. The emission of GWs from an inspiraling binary system composed of two black holes in the non-relativistic and circular orbit approximations was analyzed, and the parameters of the model were constrained by observational data. Although this work yielded interesting results, it neglected the contributions of spin-2 longitudinal polarizations. Additionally, it did not investigate the loss of angular momentum due to GW emission in the context of quadratic gravity. The present paper aims to address these two issues.
@@
The non-conservation of angular momentum in gravitational systems plays an important role both in the dynamics of the physical system that generates the GWs and in the dynamics of the GWs themselves. This is a key focus of the present paper: evaluating the effect of angular momentum transfer from the source to the GWs. This analysis requires careful consideration, as in the linear regime, quadratic actions behave like classical field theories with Lagrangians that present second-order derivatives. From this perspective, we cannot apply the known formulas for the angular momentum current usually derived in the first-order case. This will demand a proper extension, which will presented here.
@@
In the context of GWs, the term $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$, associated with the massive spin-$2$ mode, encompasses five degrees of freedom: two transverse and three longitudinal. Furthermore, this term is also responsible for introducing the Ostrogradski instability, which is evident through a negative sign in the energy density carried away by the gravitational wave. Focusing on the specific case of binary systems in the circular orbit approximation, Ref. [40] demonstrated that it is possible to consistently manage this instability by restricting the analysis to the transverse degrees of freedom. However, does this consistency extend to more general scenarios that include longitudinal polarizations and/or account for angular momentum loss? Addressing this question is the main motivation for this paper.
@@
In the present paper, we intend to generalize the analysis of the emission of GWs in quadratic gravity are presented in Ref. [40]. We will start by presenting, in Section II, the linear approximation for our quadratic model and the resulting wave equations for the perturbations. Therein, we obtain the conserved currents for our system which are related to the energy-momentum tensor and angular momentum current, both for the GWs. Next, we evaluate the energy and angular momentum carried away by the GWs. In Section III, we restrict the analysis to the quadrupole approximation and consider the propagation of GWs in a generic direction. Then, we evaluate the power radiated and consider a first example of a binary system in circular orbit. This section also includes a detailed discussion of the Ostrogradski instability and a possible solution to this issue. In Section IV, the source of GWs is taken as a rotating ellipsoid with free precession, and the role played by the quadratic terms is analyzed in comparison with the predictions of GR. At last, in Section V, we present our final comments.
@@ II Gravitational waves in quadratic gravity
In this paper, we shall carefully study the process of emission and propagation of gravitational waves in a flat background in the context of QG. We write the metric tensor, as usual, [47],
@@
and, inspired by Ref. [40], decompose $\bar{h}_{\mu\nu}$ in the form
$\bar{h}_{\mu\nu}=\tilde{h}_{\mu\nu}+\Psi_{\mu\nu}-\eta_{\mu\nu}\left(\Phi+\frac{1}{3}\Psi\right).$ (3)
@@
$\sqrt{-g}R$ $=-\frac{1}{4}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\tilde{h}^{\alpha\beta}-\frac{1}{4}\partial_{\mu}\Psi_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}$
$+\frac{3}{2}\partial_{\alpha}\Phi\partial^{\alpha}\Phi-\frac{1}{2}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}$
$\quad-\partial_{\beta}\tilde{h}^{\alpha\beta}\partial_{\alpha}\Phi-\partial_{\beta}\Psi^{\alpha\beta}\partial_{\alpha}\Phi$
$-\frac{1}{2}\partial_{\mu}\tilde{h}\partial^{\mu}\Phi+\frac{1}{2}\partial_{\alpha}\Phi\partial^{\alpha}\Psi,$
$\sqrt{-g}R^{2}$ $=$ $9\left(\square\Phi\right)^{2}-6\partial^{\mu}\partial^{\nu}\tilde{h}_{\mu\nu}\square\Phi$
$-6\partial^{\mu}\partial^{\nu}\Psi_{\mu\nu}\square\Phi$
$-3\square\tilde{h}\square\Phi+3\square\Psi\square\Phi,$
@@
$\sqrt{-g}C^{2}$ $=$ $\frac{1}{2}\square\tilde{h}_{\mu\beta}\square\tilde{h}^{\mu\beta}+\square\tilde{h}_{\mu\beta}\square\Psi^{\mu\beta}$
$+\frac{1}{2}\square\Psi_{\mu\beta}\square\Psi^{\mu\beta},$
where $\square=\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}$ and $C^{2}=C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$. Eqs. (II), (II), and (II) assume the harmonic-transverse-traceless-like gauge [40]
@@
$\mathcal{L}$ $=\frac{1}{2\kappa}\left[\frac{1}{4}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\tilde{h}^{\alpha\beta}+\frac{1}{4}\partial_{\mu}\Psi_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}\right.$ (9)
$-\frac{3}{2}\partial_{\alpha}\Phi\partial^{\alpha}\Phi+\frac{1}{2}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}$
$+\partial_{\beta}\tilde{h}^{\alpha\beta}\partial_{\alpha}\Phi+\partial_{\beta}\Psi^{\alpha\beta}\partial_{\alpha}\Phi$
@@
$+\frac{\gamma}{2\kappa}\left[-\frac{9}{2}\left(\square\Phi\right)^{2}+3\partial_{\alpha}\partial_{\beta}\tilde{h}^{\alpha\beta}\square\Phi\right.$
$+3\partial_{\alpha}\partial_{\beta}\Psi^{\alpha\beta}\square\Phi+\frac{3}{2}\square\tilde{h}\square\Phi$
$\left.-\frac{3}{2}\square\Psi\square\Phi\right]$
$+\frac{\alpha}{2\kappa}\left[\frac{1}{4}\square\tilde{h}_{\alpha\beta}\square\tilde{h}^{\alpha\beta}+\frac{1}{2}\square\tilde{h}_{\alpha\beta}\square\Psi^{\alpha\beta}\right.$
$\left.+\frac{1}{4}\square\Psi_{\mu\beta}\square\Psi^{\mu\beta}\right].$
@@
Eq. (10) is obviously of the Kein-Gordon type. The tensorial equation (11) admits a massive mode and a massless solution. These facts motivate the definitions:
@@
With that, the final form of the field equation are: [^2] The field equations in (13)–(15) are identical to those in Ref. [40] after the mapping $\alpha\rightarrow\alpha/2$. This numerical factor difference in the coupling usually appears in the context of higher-order theories in connection with the strategy adopted for the approximative scheme, i.e. whether the approximation is taken at the level of the action or at the level of the field equations. In any case, the physical content in both alternatives is the same.
@@ II.1 Conserved currents
Accordingly, infinitesimal transformations take on the form [47]
@@
The field theory formalism teaches us that symmetry transformations leaving the action invariant produce conserved quantities associated to Noether currents. These currents can be determined from the specific Lagrangian density at hand and from the functions $A_{a}^{\mu}$ e $F_{\alpha\beta,a}$ as follows.
Usually, one deals with first-order Lagrangian densities, i.e. those depending on the field and its first-order derivatives. However, quadratic gravity requires a second-order Lagrangian, viz. $\mathcal{L}=\mathcal{L}\left(\bar{h}_{\alpha\beta},\partial_{\mu}\bar{h}_{\alpha\beta},\partial_{\mu}\partial_{\nu}\bar{h}_{\alpha\beta}\right)$. As a consequence, the structural form of the current in terms of $\mathcal{L}$ must be generalized. The expression for the conserved current $j_{a}^{\mu}$ related to the second-order Lagrangian is derived in Appendix A; it reads:
$j_{a}^{\mu}=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\bar{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\bar{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)-\mathcal{L}A_{a}^{\mu}\left(x\right)$ (18)
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}h_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\bar{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)$
$+\partial_{\nu}\left\{\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\bar{h}_{\alpha\beta}\right)}\left(A_{a}^{\sigma}\partial_{\sigma}\bar{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)\right\}.$
@@
$j_{a}^{\mu}=\left(j_{a}^{\mu}\right)^{\tilde{h}}+\left(j_{a}^{\mu}\right)^{\Psi}+\left(j_{a}^{\mu}\right)^{\Phi}-\mathcal{L}A_{a}^{\mu}\left(x\right)$ (19)
@@
$\left(j_{a}^{\mu}\right)^{\tilde{h}}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\tilde{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)$ (20)
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\tilde{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right),$
$\left(j_{a}^{\mu}\right)^{\Psi}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\Psi_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right)$ (21)
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\Psi_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right),$
$\left(j_{a}^{\mu}\right)^{\Phi}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\Phi}\left(A_{a}^{\sigma}\partial_{\sigma}\Phi-F_{\alpha\beta,a}\right)$ (22)
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\Phi}\left(A_{a}^{\sigma}\partial_{\sigma}\Phi-F_{\alpha\beta,a}\right).$
@@
$\left(j_{a}^{\mu}\right)^{\tilde{h}}$ $=\frac{1}{4\kappa}\partial^{\mu}\tilde{h}^{\alpha\beta}\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)$ (23)
@@
$-\frac{1}{4\kappa}\eta^{\alpha\beta}\partial^{\mu}\Phi\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right),$
$\left(j_{a}^{\mu}\right)^{\Psi}$ $=\frac{1}{4\kappa}\partial^{\mu}\tilde{h}^{\alpha\beta}\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right)$ (24)
@@
$-\frac{1}{4\kappa}\eta^{\alpha\beta}\partial^{\mu}\Phi\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right),$

### v2

@@ I Introduction
Gravitational waves are certainly one of the most disruptive discoveries in astrophysics in recent years. Although its prediction dates back to 1916 Einstein (1916), the first evidence of its existence was only glimpsed in the 1970s by Hulse and Taylor Hulse and Taylor (1975), who observed that the system PSR 1913+16 lost energy and angular momentum in accordance with the theory of General Relativity (GR). It took another 40 years for the first direct detection of a gravitational wave (GW), due to the merger of two black holes, by the LIGO collaboration in 2015 Abbott (2016). This event marked the opening of a new window of observations, creating a new perspective for understanding astrophysical phenomena. Two years later, a major breakthrough was achieved with the observation of the merger of two neutron stars by the LIGO-VIRGO collaboration Abbott (2017). This was the first event in which both the GW and its electromagnetic counterpart were observed together, marking the beginning of the era of multimessenger astronomy.
@@
Besides allowing the comprehension of astrophysics phenomena from a broader perspective, GW is also a powerful tool for studying the gravitational interaction itself. Although GR is the standard theory for gravitation, with several successful predictions, it fails to describe galaxy rotation curves and the accelerated expansion of the Universe in the presence of ordinary matter Bertone and Hooper (2018); Caldwell and Kamionkowski (2009); Astier (2006); Riess (2004); the consistent description of the early (inflationary) Universe is hardly achieved by GR without the presence of extra degrees of freedom Starobinsky (1980); Guth (1981); Martin et al. (2014); Cuzinatto et al. (2019a); Rodrigues-da Silva et al. (2022); Bezerra-Sobrinho and Medeiros (2023); Rodrigues-da Silva and Medeiros (2023). Also, the non-renormalizability and the presence of singularities Utiyama and DeWitt (1962); Stelle (1977) do not allow us to have a self-consistent quantum version of GR with the standard (perturbative) prescription of quantization. Due to these limitations, there is almost a consensus on the scientific community that GR should not be taken as the definitive theory of gravitation. This motivated the search for new theories of gravity.
@@
The proposition of new theories of gravitation can be done by a plethora of different approaches. They include the modification of the geometry Hayashi and Shirafuji (1979); Arcos and Pereira (2004); Casana et al. (2007); Cuzinatto et al. (2021, 2008); Battista and De Falco (2021, 2022), the modification of the gravitational action Sotiriou and Faraoni (2010); Capozziello and De Laurentis (2011); Novello et al. (2013); Rinaldi et al. (2014); Pucheu et al. (2016a, b); Cuzinatto et al. (2016, 2019b); Nojiri and Odintsov (2011); Nojiri et al. (2017); Edelstein et al. (2021) (in the Riemann manifold), and many others. Usually, all these proposals share one same concern: one should obtain GR (or GR-equivalent theories) as a limiting case. With this in mind, we can understand that one of the simplest extensions to GR, besides considering the cosmological constant, consists of considering contributions to the gravitational action integral of terms that are quadratic in the curvature (this is, for instance, the case of the Starobinsky model of inflation Starobinsky (1980)). The presence of quadratic terms in the action (quadratic gravity), however, brings with them a significant modification of the dynamical equations: the field equations are of fourth order in the derivative of the metric tensor.
@@
Theories with higher-order derivatives are quite appealing in the context of quantum gravity, and some of the problems of GR may eventually be worked around – for instance, it was shown by Stelle Stelle (1977), that a renormalizable theory of gravity can be obtained for the quadratic Lagrangian. Despite being renormalizable, quadratic gravity exhibits Ostrogradsky instability, which, when quantized, gives rise to ghost-like fields Sbisà (2015). The quantization of these fields is a nontrivial task and often leads to issues with unitarity Pais and Uhlenbeck (1950). One way to address this issue is to recognize that ghosts are unstable and, as a result, do not contribute to the asymptotic spectrum of the theory Modesto and Shapiro (2016); Donoghue and Menezes (2019). Alternatively, one could define a nontrivial norm for ghost states to restore their unitarity Salvio (2019, 2021). While there is currently no consensus in the literature on how to handle the ghosts present in higher-order models, quadratic gravity and its extensions Rachwal et al. (2021); Asorey et al. (1997); Modesto and Shapiro (2016); ”Rachwal (2022) can be treated as effective theories that implement corrections to GR.
@@
The investigation of gravitational waves within the framework of quadratic gravity was particularly analyzed in Ref. Alves et al. (2023). [^1] See also Refs. de Rey Neto (2003); ”Bogdanos et al. (2010); Capozziello et al. (2017); Hölscher (2019); ”Kim et al. (2021); ”Chowdhury et al. (2023); Odintsov and Oikonomou (2022); Alves (2024) for related works. In this study, a modification to the Einstein-Hilbert action was proposed, incorporating quadratic contributions from the scalar curvature and the Weyl tensor, i.e., $R^{2}$ and $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$ terms, respectively. The GWs obtained in this work exhibited two spin-2 modes—one massive and one massless—along with a scalar (spin-0) mode. The emission of GWs from an inspiraling binary system composed of two black holes in the non-relativistic and circular orbit approximations was analyzed, and the parameters of the model were constrained by observational data. Although this work yielded interesting results, it neglected the contributions of the massive spin-2 longitudinal polarizations. Additionally, it did not investigate the loss of angular momentum due to GW emission in the context of quadratic gravity. The present paper aims to address these two issues.
@@
The non-conservation of angular momentum in gravitational systems plays an important role both in the dynamics of the physical system that generates the GWs and in the dynamics of the GWs themselves. This is a key focus of the present paper: evaluating the effect of angular momentum transfer from the source to the GWs. This analysis requires careful consideration, as in the linear regime, quadratic actions behave like classical field theories with Lagrangians that present second-order derivatives. From this perspective, we cannot apply the known formulas for the angular momentum current usually derived in the first-order case. Specifically, the relation between the energy ($E$) loss and angular momentum ($J$) loss is given by $\dot{E}=\omega\dot{J}$ for quasi-circular orbits ($\omega$ is the angular velocity and an overdot denotes a time derivative). We will see below in this paper that this relation is not exactly valid in the context of quadratic gravity. This will demand a proper extension, which will be presented here.
@@
In the context of GWs, the term $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$, associated with the massive spin-$2$ mode, encompasses five degrees of freedom: two transverse and three longitudinal (two of which are linearly dependent with respect to the remaining degrees of freedom of the metric theory due to covariance requirements–more on this below in the text). Furthermore, this term is also responsible for introducing the Ostrogradsky instability, which is evident through a negative sign in the energy density carried away by the gravitational wave. Focusing on the specific case of binary systems in the circular orbit approximation, Ref. Alves et al. (2023) demonstrated that it is possible to consistently manage this instability by restricting the analysis to the transverse degrees of freedom. However, does this consistency extend to more general scenarios that include longitudinal polarizations and/or account for angular momentum loss? Addressing this question is the main motivation for this paper.
@@
In the present paper, we intend to generalize the analysis of the emission of GWs in quadratic gravity presented in Ref. Alves et al. (2023). We will start by presenting, in Section II, the linear approximation for our quadratic model and the resulting wave equations for the perturbations. Therein, we obtain the conserved currents for our system which are related to the energy-momentum tensor and angular momentum current, both for the GWs. Next, we evaluate the energy and angular momentum carried away by the GWs. In Section III, we restrict the analysis to the quadrupole approximation and consider the propagation of GWs in a generic direction. Then, we evaluate the power radiated and consider a first example of a binary system in circular orbit. This section also includes a detailed discussion of the Ostrogradsky instability and a possible solution to this issue. In Section IV, the source of GWs is taken as a rotating ellipsoid with free precession, and the role played by the quadratic terms is analyzed in comparison with the predictions of GR. At last, in Section V, we present our final comments.
@@ II Gravitational waves in quadratic gravity
In this paper, we shall carefully study the process of emission and propagation of gravitational waves in a flat background in the context of QG. We write the metric tensor, as usual, Maggiore (2007),
@@
and, inspired by Ref. Alves et al. (2023), decompose $\bar{h}_{\mu\nu}$ in the form
$\bar{h}_{\mu\nu}=\tilde{h}_{\mu\nu}+\Psi_{\mu\nu}-\eta_{\mu\nu}\left(\Phi+\frac{1}{6}\Psi\right).$ (3)
@@
$\sqrt{-g}R$ $=-\frac{1}{4}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\tilde{h}^{\alpha\beta}-\frac{1}{4}\partial_{\mu}\Psi_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}+\frac{3}{2}\partial_{\alpha}\Phi\partial^{\alpha}\Phi$
$-\frac{1}{2}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}-\partial_{\beta}\tilde{h}^{\alpha\beta}\partial_{\alpha}\Phi-\partial_{\beta}\Psi^{\alpha\beta}\partial_{\alpha}\Phi$
$-\frac{1}{2}\partial_{\mu}\tilde{h}\partial^{\mu}\Phi+\frac{1}{2}\partial_{\alpha}\Phi\partial^{\alpha}\Psi,$ (4)
$\sqrt{-g}R^{2}$ $=9\left(\square\Phi\right)^{2}-6\partial^{\mu}\partial^{\nu}\tilde{h}_{\mu\nu}\square\Phi-6\partial^{\mu}\partial^{\nu}\Psi_{\mu\nu}\square\Phi$
$-3\square\tilde{h}\square\Phi+3\square\Psi\square\Phi,$ (5)
@@
$\sqrt{-g}C^{2}=\frac{1}{2}\square\tilde{h}_{\mu\beta}\square\tilde{h}^{\mu\beta}+\square\tilde{h}_{\mu\beta}\square\Psi^{\mu\beta}+\frac{1}{2}\square\Psi_{\mu\beta}\square\Psi^{\mu\beta},$ (6)
where $\square=\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}$ and $C^{2}=C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$. Eqs. (4), (5), and (6) assume the harmonic-transverse-traceless-like gauge Alves et al. (2023)
@@
$\mathcal{L}$ $=\frac{1}{2\kappa}\left[\frac{1}{4}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\tilde{h}^{\alpha\beta}+\frac{1}{4}\partial_{\mu}\Psi_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}-\frac{3}{2}\partial_{\alpha}\Phi\partial^{\alpha}\Phi\right.$
$\qquad+\frac{1}{2}\partial_{\mu}\tilde{h}_{\alpha\beta}\partial^{\mu}\Psi^{\alpha\beta}+\partial_{\beta}\tilde{h}^{\alpha\beta}\partial_{\alpha}\Phi+\partial_{\beta}\Psi^{\alpha\beta}\partial_{\alpha}\Phi$
@@
$+\frac{\gamma}{2\kappa}\left[-\frac{9}{2}\left(\square\Phi\right)^{2}+3\partial_{\alpha}\partial_{\beta}\tilde{h}^{\alpha\beta}\square\Phi+3\partial_{\alpha}\partial_{\beta}\Psi^{\alpha\beta}\square\Phi\right.$
$\qquad\left.+\frac{3}{2}\square\tilde{h}\square\Phi-\frac{3}{2}\square\Psi\square\Phi\right]$
$+\frac{\alpha}{2\kappa}\left[\frac{1}{4}\square\tilde{h}_{\alpha\beta}\square\tilde{h}^{\alpha\beta}+\frac{1}{2}\square\tilde{h}_{\alpha\beta}\square\Psi^{\alpha\beta}+\frac{1}{4}\square\Psi_{\mu\beta}\square\Psi^{\mu\beta}\right].$ (9)
@@
Eq. (10) is obviously of the Klein-Gordon type. The tensorial equation (11) admits a massive mode and a massless solution. These facts motivate the definitions:
@@
With that, the final form of the field equation are:222 The field equations in (13)–(15) are identical to those in Ref. Alves et al. (2023) after the mapping $\alpha\rightarrow\alpha/2$. This numerical difference in the coupling factor appeared due to a small error in Ref. Alves et al. (2023). In equation (2) of this reference, we should have $2\alpha$ instead of $\alpha$. This correction was made in the current paper and led to the new definition of $m_{\Psi}^{2}$ without the factor $2$ that appears in Ref. Alves et al. (2023). This same correction leads to a change from $1/3$ of Eq. (2) in Ref. Alves et al. (2023) to $1/6$ in Eq. (3) of the current paper.
@@ II.1 Conserved currents
Accordingly, infinitesimal transformations take on the form Maggiore (2007)
@@
The field theory formalism teaches us that symmetry transformations leaving the action invariant produce conserved quantities associated with Noether currents. These currents can be determined from the specific Lagrangian density at hand and from the functions $A_{a}^{\mu}$ e $F_{\alpha\beta,a}$ as follows. Usually, one deals with first-order Lagrangian densities, i.e. those depending on the the field and its first-order derivatives. However, quadratic gravity requires a second-order Lagrangian, viz. $\mathcal{L}=\mathcal{L}\left(\bar{h}_{\alpha\beta},\partial_{\mu}\bar{h}_{\alpha\beta},\partial_{\mu}\partial_{\nu}\bar{h}_{\alpha\beta}\right)$. As a consequence, the structural form of the current in terms of $\mathcal{L}$ must be generalized. The expression for the conserved current $j_{a}^{\mu}$ related to the second-order Lagrangian is derived in Appendix A; it reads:
$j_{a}^{\mu}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\bar{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\bar{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)-A_{a}^{\mu}\mathcal{L}$
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\bar{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\bar{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)$
$+\partial_{\nu}\left\{\frac{\partial\mathcal{L}}{\partial\left(\partial_{\mu}\partial_{\nu}\bar{h}_{\alpha\beta}\right)}\left(A_{a}^{\sigma}\partial_{\sigma}\bar{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)\right\}.$ (18)
@@
$j_{a}^{\mu}=\left(j_{a}^{\mu}\right)^{\tilde{h}}+\left(j_{a}^{\mu}\right)^{\Psi}+\left(j_{a}^{\mu}\right)^{\Phi}-A_{a}^{\mu}\mathcal{L}$ (19)
@@
$\left(j_{a}^{\mu}\right)^{\tilde{h}}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\tilde{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)$
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\tilde{h}_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right),$ (20)
$\left(j_{a}^{\mu}\right)^{\Psi}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\Psi_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right)$
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\Psi_{\alpha\beta}}\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right),$ (21)
$\left(j_{a}^{\mu}\right)^{\Phi}$ $=\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\Phi}\left(A_{a}^{\sigma}\partial_{\sigma}\Phi-F_{\alpha\beta,a}\right)$
$-2\partial_{\nu}\frac{\partial\mathcal{L}}{\partial\partial_{\mu}\partial_{\nu}\Phi}\left(A_{a}^{\sigma}\partial_{\sigma}\Phi-F_{\alpha\beta,a}\right).$ (22)
@@
$\left(j_{a}^{\mu}\right)^{\tilde{h}}$ $=\frac{1}{4\kappa}\partial^{\mu}\tilde{h}^{\alpha\beta}\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right)$
@@
$-\frac{1}{4\kappa}\eta^{\alpha\beta}\partial^{\mu}\Phi\left(A_{a}^{\sigma}\partial_{\sigma}\tilde{h}_{\alpha\beta}-F_{\alpha\beta,a}\right),$ (23)
$\left(j_{a}^{\mu}\right)^{\Psi}$ $=\frac{1}{4\kappa}\partial^{\mu}\tilde{h}^{\alpha\beta}\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right)$
@@
$-\frac{1}{4\kappa}\eta^{\alpha\beta}\partial^{\mu}\Phi\left(A_{a}^{\sigma}\partial_{\sigma}\Psi_{\alpha\beta}-F_{\alpha\beta,a}\right),$ (24)